OpenStudy (anonymous):

what is the solution to the equation 1/2x + 3/4 = 1/4 for x? c; Thanks for helping~

OpenStudy (johnweldon1993):

Well we want to isolate 1/2x over there on the left side....so we should subtract 3/4 from both sides....3/4 - 3/4 = 0 right? so this will leave 1/2x alone 1/2x = (1/4) - (3/4) first...what is 1/4 - 3/4?

OpenStudy (johnweldon1993):

close.... 1/4 - 3/4 = (1 - 3)/4 = -2/4 -2/4 simplified would be -1/2 right? (dividing both sides by 2) so we have 1/2x = -1/2 so what does x = ...?

OpenStudy (anonymous):

x = -1? >_>

OpenStudy (johnweldon1993):

Haha what's with the face..that would be the correct answer....Did you not understand what we did there?

OpenStudy (anonymous):

Yes I did! Thank you c:

OpenStudy (johnweldon1993):

Lol okay good! and anytime :)